OpenStudy (anonymous):

Solve for X: -5 |x + 1| = 10 a. x = 0 b. x = -3 and x = 1 c. x = -1 and x = 3 d. no solution

jimthompson5910 (jim_thompson5910):

what happens when you divide both sides by -5?

OpenStudy (anonymous):

I got -2 @jim_thompson5910

OpenStudy (freckles):

Can distance be negative?

OpenStudy (freckles):

Does anything have a distance of -2 from 0?

jimthompson5910 (jim_thompson5910):

correct, you should get -2 on the right side. Do you understand what freckles is asking? or no?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

absolute value represents distance for example, |-3| = 3 means that -3 is three units away from 0 if we said |x| = 2, then the answers are x = -2 or x = 2 because each number is two units away from 0. Notice how the distance is never negative

jimthompson5910 (jim_thompson5910):

so saying |x + 1| = -2 is impossible therefore, there are no solutions to this equation or the original equation

jimthompson5910 (jim_thompson5910):

there are no values of x that will make |x + 1| = -2 true there are no values of x that will make -5|x + 1| = 10 true

OpenStudy (anonymous):

oh okay, thank you so much! (:

jimthompson5910 (jim_thompson5910):

you're welcome